# On the multiplier spectrum of polynomials

### v1

@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
The case $l=1$ follows from Proposition 3.2. Assume $l\geq 2$ and consider a decomposition $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ into indecomposable rational maps. By considering the totally ramified fixed point $\infty$, we may assume $F_{r},\dots,F_{1}$ are polynomials. By [ZM08, Theorem 1.3] and Proposition 3.2, we have $r=l$, $m=\deg(F_{l})=\cdots=\deg(F_{1})$, and ${\rm Mon}(F_{j})\cong S_{m}\cong{\rm Mon}(F)$ for all $1\leq j\leq l$. The tuple $(m=\deg(F_{l}),\dots,m=\deg(F_{1}))$ and the isomorphic type $S_{m}$ of ${\rm Mon}(F_{i})$ are invariants of complete decompositions of $F^{\circ l}$.
Suppose the decomposition $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ is not equivalent to $F^{\circ l}=F\circ\cdots\circ F$. Then by Ritt’s theorem on polynomial decomposition [FG22, Theorem 3.35] (see [Rit22, ZM08]), there exist $j\in\{1,\dots,l\}$ and linear polynomials $\sigma_{1}(z),\sigma_{2}(z)\in\mathbb{C}[z]$ such that $\sigma_{1}\circ F_{j}\circ\sigma_{2}(z)=z^{m}$. (Note that $\gcd(m,m)=m\neq 1$, so we need not consider the Chebyshev polynomial $T_{m}(z)$; see [FG22, §3.5.1 (M3)].) Then $S_{m}\cong{\rm Mon}(F_{i})\cong{\rm Mon}(\sigma_{1}\circ F_{j}\circ\sigma_{2})={\rm Mon}(z^{m})\cong\mathbb{Z}/m\mathbb{Z}$ (see [ZM08, Lemma 3.6]), contradicting $m\geq 3$. ∎
@@
In Theorem 3.7, we only require $m\geq 3$. The proof relies on Ritt’s theory on polynomial decompositions. For $m\geq 5$, an alternative proof following [Pak25a] is as follows:
We proceed by induction on $l$. The case $l=1$ follows from Proposition 3.2. Assume $l\geq 2$ and the result holds for l1l-1. Consider a decomposition $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ into indecomposable rational maps, with $F_{r},\dots,F_{1}$ polynomials.
Since $m\geq 5$, Theorem 3.5 implies that the algebraic curve $F(x)-F_{r}(y)=0$ is reducible (otherwise $F\circ X=F_{r}\circ Y$ has no solutions). By Proposition 3.2 and [Pak25a, Theorem 2.7], either $F_{r}=F\circ\mu$ for some $\mu\in{\rm PGL}_{2}(\mathbb{C})$, or $\deg(F_{r})=\binom{m}{k}$ for some 1<k<m11<k<m-1. If $F_{r}=F\circ\mu$, then $F^{\circ(l-1)}=(\mu^{-1}\circ F_{r-1})\circ\cdots F_{1}$ by Theorem 3.6, and the conclusion follows by induction. If $\deg(F_{r})=\binom{m}{k}$ for some 1<k<m11<k<m-1, then by [Pak25a, Theorem 2.9], since $m\geq 4$, there exists a prime $p$ dividing $\binom{m}{k}=\deg(F_{r})$ such that $p\nmid m$. But $p$ divides $m^{l}=\deg(F^{\circ l})$ since $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$, which is a contradiction. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$, and let $G_{1},\dots,G_{r}\in\mathbb{C}(z)$ be rational maps of degree $\geq 2$ such that $F^{\circ l}=G_{r}\circ\cdots\circ G_{1}$ for some $l\in\mathbb{Z}_{>0}$. Then there exist $s_{1},\dots,s_{r}\in\mathbb{Z}_{>0}$ and Möbius transformations $\nu_{1},\dots,\nu_{r-1}\in\mathbb{C}(z)$ of degree $1$ such that
$G_{r}=F^{\circ s_{r}}\circ\nu_{r-1},\quad G_{j}=\nu_{j}^{-1}\circ F^{\circ s_{j}}\circ\nu_{j-1}\ (1<j<r),\quad G_{1}=\nu_{1}^{-1}\circ F^{\circ s_{1}}.$
The $s_{1},\dots,s_{r}$ are uniquely determined by $m=\deg(F)$ and the degrees of the $G_{j}$’s, with $\sum_{j=1}^{r}s_{j}=l$. If $G_{1},\dots,G_{r}$ are polynomials, then $\nu_{1},\dots,\nu_{r-1}$ can be chosen as linear polynomials.
@@
Decompose each $G_{i}$ ($1\leq i\leq r$) into indecomposable rational functions and apply Theorem 3.7. If $G_{1},\dots,G_{r}$ are polynomials, then using the totally ramified fixed point $\infty$, we see that $\nu_{1},\dots,\nu_{r-1}$ are linear polynomials. ∎
We now recall certain (semi)groups (under composition) associated with a rational map $F(z)\in\mathbb{C}(z)$ of degree $m\geq 2$, following [Pak25a].
@@
${\rm Aut}(F):=C(F)\cap{\rm PGL}_{2}(\mathbb{C})\quad\text{and}\quad{\rm Aut}_{\infty}(F):=C_{\infty}(F)\cap{\rm PGL}_{2}(\mathbb{C}).$
Both ${\rm Aut}(F)$ and ${\rm Aut}_{\infty}(F)$ are groups. Note that the semigroup $\left\langle{\rm Aut}_{\infty}(F),F\right\rangle$ generated by ${\rm Aut}_{\infty}(F)\cup\{F\}$ is contained in $C_{\infty}(F)$.
For $g\in\mathbb{C}(z)$ of degree $\geq 2$, let $\mu_{g}$ denote the measure of maximal entropy of $g$ (see [Lyu83, Mañ83]). Define
$E_{0}(F):=\{\sigma\in{\rm PGL}_{2}(\mathbb{C})\mid\sigma_{*}\mu_{F}=\mu_{F}\}$
@@
Define $G(F)$ as the subgroup of ${\rm PGL}_{2}(\mathbb{C})$ given by
$\{\sigma\in{\rm PGL}_{2}(\mathbb{C}):\exists\tau\in{\rm PGL}_{2}(\mathbb{C}),\,F\circ\sigma=\tau\circ F\}.$
Define $G_{0}(F)$ as the maximal subgroup of ${\rm PGL}_{2}(\mathbb{C})$ such that for every $\sigma\in G_{0}(F)$, there exists $\tau\in G_{0}(F)$ with $F\circ\sigma=\tau\circ F$. Clearly, $G_{0}(F)\subseteq G(F)$.
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then $G_{0}(F)$ is finite. Define $\gamma:G(F)\to{\rm PGL}_{2}(\mathbb{C})$ by $\sigma\mapsto\tau_{\sigma}$, where $\tau_{\sigma}$ is the unique Möbius transformation such that $F\circ\sigma=\tau_{\sigma}\circ F$. Then the restriction $\gamma:G_{0}(F)\to\gamma(G_{0}(F))$ is a group automorphism of $G_{0}(F)$.
@@
By [Pak20b, §4], the group $G(F)$ is finite of order bounded in terms of $m=\deg(F)$, unless $(\alpha\circ F\circ\beta)(z)=z^{m}$ for some $\alpha,\beta\in{\rm PGL}_{2}(\mathbb{C})$. Since $F$ is pre-simple, for any Möbius transformations $\alpha$ and $\beta$, the map $\alpha\circ F\circ\beta$ is pre-simple and has exactly $m$ critical values in $\widehat{\mathbb{C}}$. But $z^{m}$ has only $2(<m)$ critical values in $\widehat{\mathbb{C}}$. Hence, $G(F)$ is finite, and so is its subgroup $G_{0}(F)$. By definition, $\gamma(G_{0}(F))\subseteq G_{0}(F)$. By Theorem 3.6, $\gamma$ is injective for $m\geq 4$. Thus, $\gamma:G_{0}(F)\to G_{0}(F)$ is a bijection. It is easy to check that $\gamma$ is a group homomorphism, hence an automorphism. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$, and let $s=\#G_{0}(F)\in\mathbb{Z}_{>0}$. Then $G_{0}(F)\subseteq{\rm Aut}(F^{\circ s})$.
@@
By Lemma 3.10, $\gamma:G_{0}(F)\to\gamma(G_{0}(F))$ is a group automorphism of the finite group $G_{0}(F)$ of order $s$, so $\gamma^{\circ s}={\rm Id}_{G_{0}(F)}$. For every $\sigma\in G_{0}(F)$, we have
@@
hence $G_{0}(F)\subseteq{\rm Aut}(F^{\circ s})$. ∎
@@
Suppose $F$ is a pre-simple polynomial of degree $m\geq 4$ that is exceptional. Since $F$ is a polynomial, it is not a Lattès map. Then $F$ is of monomial type, i.e., conjugate to $z^{\pm m}$ or $\pm T_{m}(z)$, where $T_{m}$ is the Chebyshev polynomial of degree $m$. As $F$ is a polynomial, it cannot be conjugate to $z^{-m}$. Hence, $F$ is conjugate to $z^{m}$, $T_{m}(z)$, or $T_{m}(z)$ by a linear polynomial. Then $z^{m}$ or $T_{m}(z)$ must be pre-simple. But $z^{m}$ is not pre-simple for $m\geq 3$. The Chebyshev polynomial $T_{m}$ satisfies $(4-z^{2})F^{\prime}(z)^{2}=m^{2}(4-(F(z))^{2})$, so it has at most $2$ distinct critical values ($\pm 2$) in $\mathbb{C}$. Since 2<m12<m-1, $T_{m}$ is not pre-simple and we get a contradiction. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then
$G_{0}(F)={\rm Aut}(F^{\circ s})={\rm Aut}_{\infty}(F)=E_{0}(F),\quad C_{\infty}(F)=\left\langle{\rm Aut}_{\infty}(F),F\right\rangle=E(F),$
where $s=\#G_{0}(F)\in\mathbb{Z}_{>0}$.
@@
We have $G_{0}(F)\subseteq{\rm Aut}(F^{\circ s})\subseteq{\rm Aut}_{\infty}(F)\subseteq E_{0}(F)$, where the first inclusion is by Corollary 3.11, the second by definition, and the third by [Pak25a, (31)].
To show $E_{0}(F)\subseteq G_{0}(F)$, let $\sigma\in E_{0}(F)$. Then $F\circ\sigma\in E(F)$. By Lemma 3.12, $F$ is non-exceptional. By [Lev90], there exist integers $k_{1},l>0$ and $k_{2}\geq 0$ with $k_{1}=k_{2}+l$ such that $F^{\circ k_{1}}=F^{\circ k_{2}}\circ(F\circ\sigma)^{\circ l}$. Applying Theorem 3.6 recursively, we get $F^{\circ l}=F^{\circ(k_{1}-k_{2})}=(F\circ\sigma)^{\circ l}$. By Theorem 3.7, there exist $\mu_{1},\dots,\mu_{l-1}\in{\rm PGL}_{2}(\mathbb{C})$ such that
$F\circ\sigma=F\circ\mu_{l-1},\quad F\circ\sigma=\mu_{i}^{-1}\circ F\circ\mu_{i-1}\ (1<i<l),\quad F\circ\sigma=\mu_{1}^{-1}\circ F.$
Set $\mu_{1}(z)=z$ when $l=1$. Then $\nu:=\mu_{1}^{-1}$ satisfies $F\circ\sigma=\nu\circ F$. We claim $\nu\in E_{0}(F)$. For $l=1$, this is clear. For $l\geq 2$, note that $F^{\circ l}=(F\circ\sigma)^{\circ l}=(\nu\circ F)^{\circ l}$. By [Pak25a, Lemma 3.3], $\nu\in{\rm Aut}(F^{\circ l})\subseteq{\rm Aut}_{\infty}(F)\subseteq E_{0}(F)$. Hence, $E_{0}(F)\subseteq G_{0}(F)$.
Thus, $G_{0}(F)={\rm Aut}(F^{\circ s})={\rm Aut}_{\infty}(F)=E_{0}(F)$.

### v2

@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
Let $\tau\in\Sigma(f^{\circ n})$ for some integer $n\in\mathbb{Z}_{\geq 0}$. Then $\tau$ is a polynomial of degree one by (3.2). Applying Theorem 3.7 recursively to the equation $f^{\circ n}\circ\tau=f^{\circ n}$, we see that $\tau(z)=z$. Therefore, $\Sigma_{\infty}(f)=1$ is the trivial group and so is $\Sigma(f)$. ∎
@@
Note that an iterate $f^{\circ n}$ of a pre-simple polynomial $f\in\mathbb{C}[z]$ may not be pre-simple when $n\geq 2$; so Corollary 3.9 is not a direct result of Lemma 3.6.
By Ritt’s theory of polynomial decompositions, we see that self-compositions of an indecomposable polynomial behave well under decomposition:
Proposition 3.11.
Let $F\in\mathbb{C}[z]$ be a polynomial of degree $m\geq 2$. Suppose that $F$ is indecomposable (for example, when $F$ is pre-simple (Proposition 3.2) or $m$ is prime). Then for every $l\in\mathbb{Z}_{>0}$, every complete decomposition of $F^{\circ l}$ is equivalent to the complete decomposition $F^{\circ l}=F\circ\cdots\circ F$.
Proof.
The case $l=1$ is trivial. Assume $l\geq 2$ and let $F^{\circ l}=F_{1}\circ\cdots\circ F_{r}$ be a complete decomposition of $F^{\circ l}$. Since $F^{\circ l}=F\circ\cdots\circ F$ is a complete decomposition of $F^{\circ l}$ of length $l$, we have $r=l$ by [FG22, Theorem 3.35]. Up to equivalent (complete) decompositions, the totally invariant fixed point $\infty$ implies that we may assume that $F_{1},\dots,F_{l}$ are all polynomials.
Note that $n:=\gcd(\deg(F_{1}),m)\geq 2$ because $\deg(F_{1})\mid m^{l}=\deg(F^{\circ l})$. By [FG22, Theorem 3.54], there exist polynomials $G,\sigma,\mu\in\mathbb{C}[z]\setminus\mathbb{C}$ such that
$F=G\circ\sigma,\quad F_{1}=G\circ\mu,\quad\text{and}\quad\sigma\circ F^{\circ(l-1)}=\mu\circ F_{2}\circ\cdots\circ F_{l}$
with $\deg(G)=n>1$. Since both $F$ and $F_{1}$ are indecomposable, we conclude that
$n=\deg(G)=\deg(F_{1})=\deg(F)=m\quad\text{and}\quad\deg(\sigma)=\deg(\mu)=1.$
Set $\tau=\sigma^{-1}\circ\mu\in\mathrm{Aff}(\mathbb{C})$. Then we obtain
$F_{1}=F\circ\tau\quad\text{and}\quad F^{\circ(l-1)}=(\tau\circ F_{2})\circ F_{3}\circ\cdots\circ F_{l}.$
An inductive argument then finishes the proof. ∎
@@
Let $F\in\mathbb{C}[z]$ be a polynomial of degree $m\geq 2$. Suppose that $F$ is indecomposable (for example, when $F$ is pre-simple or $m$ is prime). Let $G_{1},\dots,G_{r}\in\mathbb{C}(z)$ be rational maps of degree $\geq 2$ such that $F^{\circ l}=G_{1}\circ\cdots\circ G_{r}$, where $r,l\in\mathbb{Z}_{>0}$. Then there exist $\nu_{1},\dots,\nu_{r-1}\in\mathrm{PGL}_{2}(\mathbb{C})$ such that
$G_{1}=F^{\circ s_{1}}\circ\nu_{1},\quad G_{j}=\nu_{j-1}^{-1}\circ F^{\circ s_{j}}\circ\nu_{j}\ (1<j<r),\quad G_{r}=\nu_{r-1}^{-1}\circ F^{\circ s_{r}},$
where $s_{i}:=\log_{m}\deg(G_{i})$ is a positive integer for $1\leq i\leq r$. If $G_{1},\dots,G_{r}$ are polynomials, then we can choose $\nu_{1},\dots,\nu_{r-1}\in\mathrm{Aff}(\mathbb{C})$.
@@
For every $1\leq i\leq r$, choose a complete decomposition $G_{i}=G_{i,1}\circ\cdots\circ G_{i,k_{i}}$ of $G_{i}$. By connecting these decompositions, we obtain a complete decomposition
(3.3) $F^{\circ l}=G_{1,1}\circ\cdots\circ G_{1,k_{1}}\circ\cdots\circ G_{r,1}\circ\cdots\circ G_{r,k_{r}}$
of $F^{\circ l}$. The conclusion then follows by applying Proposition 3.11 to the decomposition (3.3). ∎
Let $F\in\mathbb{C}(z)$ be a rational map of degree $m\geq 2$. Apart from $\Sigma(F)$ and $\Sigma_{\infty}(F)$, we recall some other (semi)groups (under composition) associated with $F$ of degree $m\geq 2$, following [Pak25a].
@@
$\mathrm{Aut}(F):=C(F)\cap\mathrm{PGL}_{2}(\mathbb{C})\quad\text{and}\quad\mathrm{Aut}_{\infty}(F):=C_{\infty}(F)\cap\mathrm{PGL}_{2}(\mathbb{C}).$
Both $\mathrm{Aut}(F)$ and $\mathrm{Aut}_{\infty}(F)$ are groups. Note that the semigroup $\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle$ generated by $\mathrm{Aut}_{\infty}(F)\cup\{F\}$ is contained in $C_{\infty}(F)$.
For $g\in\mathbb{C}(z)$ of degree $\geq 2$, let $\mu_{g}$ be the measure of maximal entropy of $g$ on $\mathbb{P}^{1}(\mathbb{C})$ (see [Lyu83, Mañ83]). Define
$E_{0}(F):=\{\sigma\in\mathrm{PGL}_{2}(\mathbb{C})\mid\sigma_{*}\mu_{F}=\mu_{F}\}$
@@
Define $G(F)$ as the subgroup of $\mathrm{PGL}_{2}(\mathbb{C})$ given by
$\{\sigma\in\mathrm{PGL}_{2}(\mathbb{C})\mid\exists\tau\in\mathrm{PGL}_{2}(\mathbb{C}),\,F\circ\sigma=\tau\circ F\}.$
Note that for every $\sigma\in G(F)$, there is a unique $\tau_{\sigma}\in\mathrm{PGL}_{2}(\mathbb{C})$ such that
(3.4) $F\circ\sigma=\tau_{\sigma}\circ F.$
The group $G(F)$ is called the extended symmetry group of $F$. Define $G_{0}(F)$ as the maximal subgroup of $\mathrm{PGL}_{2}(\mathbb{C})$ such that for every $\sigma\in G_{0}(F)$, there exists $\tau\in G_{0}(F)$ with $F\circ\sigma=\tau\circ F$. From the definitions, it is clear that
$\left\langle\mathrm{Aut}(F)\cup\Sigma(F)\right\rangle\subseteq G_{0}(F)\subseteq G(F).$
Suppose that $F$ is a polynomial. Then these semigroups associated with $F$ are expected to “relate to” polynomials. The inclusion $\Sigma_{\infty}(F)\subseteq\mathrm{Aff}(\mathbb{C})$ (see (3.2)) and the following two lemmas illustrate such an idea.
@@
Let $F\in\mathbb{C}[z]$ be a polynomial of degree $m\geq 2$ that is not conjugate to $z^{m}$. Then $C_{\infty}(F)\subseteq\mathbb{C}[z]\setminus\mathbb{C}$.
@@
We work over the algebraically closed field $\mathbb{C}$ of characteristic zero. For a polynomial $F\in\mathbb{C}[z]$ of degree $m\geq 2$, it is well-known that $F$ is conjugate to $z^{m}$ if and only if for some (hence every) $k\in\mathbb{Z}_{>0}$, the iterate $F^{\circ k}$ is conjugate to $z^{m^{k}}$; see [Mil06b, Zdu90]. Then the conclusion follows immediately from [Sil07, Theorem 6.80]. ∎
Lemma 3.14.
Let $F\in\mathbb{C}[z]$ be a polynomial of degree $m\geq 2$ that is non-exceptional. Then $E(F)\subseteq\mathbb{C}[z]\setminus\mathbb{C}$.
Proof.
We first show that $E(F)\setminus E_{0}(F)\subseteq\mathbb{C}[z]\setminus\mathbb{C}$. Let $g\in E(F)\setminus E_{0}(F)$. Since $F$ is non-exceptional, by [Lev90], there exist integers $k_{1},l>0$ and $k_{2}\geq 0$ such that $F^{\circ k_{1}}=F^{\circ k_{2}}\circ g^{\circ l}$. Note that $(g^{\circ l})^{-1}(\infty)=\{\infty\}$, hence $g^{\circ l}$ is a polynomial. Since $\mu_{g^{\circ l}}=\mu_{g}=\mu_{F}$ and $F$ is non-exceptional, $g^{\circ l}$ is also non-exceptional by [FG22, Corollary 2.17]. In particular, $g^{\circ l}$ is not conjugate to $z^{n^{l}}$, where $n:=\deg(g)\geq 2$. By Lemma 3.13, $g\in C(g^{\circ l})$ is a polynomial.
Let $\sigma\in E_{0}(F)$. Then the rational map $\tilde{F}=\sigma\circ F\circ\sigma^{-1}\in\mathbb{C}(z)$ satisfies $\mu_{\tilde{F}}=\sigma_{*}\mu_{F}=\mu_{F}$, so $\tilde{F}\in E(F)\setminus E_{0}(F)\subseteq\mathbb{C}[z]$ is a polynomial. As $\infty$ is the unique totally invariant fixed point of the non-exceptional polynomial $F$ (and of $\tilde{F}$), we have $\sigma(\infty)=\infty$. We conclude that $\sigma\in\mathrm{Aff}(\mathbb{C})$. ∎
Remark 3.15.
If $F\in\mathbb{C}[z]$ (of degree $m\geq 2$) is exceptional, then $E(F)\nsubseteq\mathbb{C}[z]\setminus\mathbb{C}$. See [Mil06a, § 7].
Lemma 3.16.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$. Define $\gamma=\gamma_{F}:G(F)\to\mathrm{PGL}_{2}(\mathbb{C}),\sigma\mapsto\tau_{\sigma}$ where $\tau_{\sigma}$ is as in (3.4). Then $G(F)$ is finite, and $\gamma$ restricts to a group automorphism of $G_{0}(F)$.
Proof.
By [Pak20b, Theorem 4.2], the group $G(F)$ is finite of order bounded in terms of $m=\deg(F)$, unless $(\alpha\circ F\circ\beta)(z)=z^{m}$ for some $\alpha,\beta\in\mathrm{PGL}_{2}(\mathbb{C})$. Since $F$ is pre-simple, for any Möbius transformations $\alpha$ and $\beta$, the map $\alpha\circ F\circ\beta$ has exactly $m$ critical values in $\widehat{\mathbb{C}}$. But $z^{m}$ has only $2(<m)$ critical values in $\widehat{\mathbb{C}}$. Hence, $G(F)$ is finite, and so is its subgroup $G_{0}(F)$. By definition, $\gamma(G_{0}(F))\subseteq G_{0}(F)$. By Theorem 3.7, $\Sigma(F)=1$, which implies that $\gamma$ is injective. Therefore, the injective self-map $\gamma:G_{0}(F)\to G_{0}(F)$ on the finite set $G_{0}(F)$ is a bijection. It is easy to check that $\gamma$ is a group homomorphism, hence an automorphism. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$. Let $s=\#\mathrm{Aut}(G_{0}(F))\in\mathbb{Z}_{>0}$, where $\mathrm{Aut}(G_{0}(F))$ is the automorphism group of the finite group $G_{0}(F)$. Then $G_{0}(F)\subseteq\mathrm{Aut}(F^{\circ s})$.
@@
By Lemma 3.16, $\gamma:G_{0}(F)\to G_{0}(F)$ is an automorphism of the finite group $G_{0}(F)$. Since $s=\#\mathrm{Aut}(G_{0}(F))$, $\gamma^{\circ s}=\mathrm{Id}_{G_{0}(F)}$ in $\mathrm{Aut}(G_{0}(F))$. For every $\sigma\in G_{0}(F)$, we have
@@
hence $G_{0}(F)\subseteq\mathrm{Aut}(F^{\circ s})$. ∎
@@
Suppose $F$ is a pre-simple polynomial of degree $m\geq 4$ that is exceptional. Since $F$ is a polynomial, it is not a Lattès map. Hence $F$ is of monomial type; so $F$ is affinely conjugate to $z^{\pm m}$ or $\pm T_{m}$, where $T_{m}\in\mathbb{C}[z]$ is the Chebyshev polynomial of degree $m$. As $F$ is a polynomial, it cannot be conjugate to $z^{-m}$. Hence, $F$ is affinely conjugate to $z^{m}$ or $\pm T_{m}$. Clearly, $z^{m}$ is not pre-simple for $m\geq 3$. It is well-known that $\pm T_{m}$ are polynomial solutions to the differential equation $(4-z^{2})G^{\prime}(z)^{2}=m^{2}(4-G(z)^{2})$ (see [Sil07, Lemma 6.10]); so the only choices of finite critical values of $\pm T_{m}$ are $\pm 2$. Since 2<m12<m-1, $\pm T_{m}$ is not pre-simple. Hence we get a contradiction to the fact that pre-simplicity is preserved by affine conjugacy. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Set $s=\#\mathrm{Aut}(G_{0}(F))$, which is a positive integer. Then:
  • $G_{0}(F)=\mathrm{Aut}(F^{\circ s})=\mathrm{Aut}_{\infty}(F)=E_{0}(F)\subseteq\mathrm{Aff}(\mathbb{C})$;
  • $C_{\infty}(F)=E(F)=\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle\subseteq\mathbb{C}[z]$.
@@
By Lemma 3.18, $F$ is non-exceptional. Then we have $E_{0}(F)\subseteq\mathrm{Aff}(\mathbb{C})$ and $E(F)\subseteq\mathbb{C}[z]$ by Lemma 3.14.
Note that $G_{0}(F)\subseteq\mathrm{Aut}(F^{\circ s})\subseteq\mathrm{Aut}_{\infty}(F)\subseteq E_{0}(F)$, where the first inclusion is by Corollary 3.17, the second by definition, and the third by [Pak25a, (29)].
We show that $E_{0}(F)\subseteq G_{0}(F)$. Let $\sigma\in E_{0}(F)\subseteq\mathrm{Aff}(\mathbb{C})$. Then $F\circ\sigma\in E(F)$. By [Lev90], there exist integers $k_{1},l>0$ and $k_{2}\geq 0$ with $k_{1}=k_{2}+l$ such that $F^{\circ k_{1}}=F^{\circ k_{2}}\circ(F\circ\sigma)^{\circ l}$. Applying Theorem 3.7 recursively, we get $F^{\circ l}=F^{\circ(k_{1}-k_{2})}=(F\circ\sigma)^{\circ l}$. By Theorem 3.11, there exist $\mu_{1},\dots,\mu_{l-1}\in\mathrm{Aff}(\mathbb{C})$ such that
$F\circ\sigma=F\circ\mu_{1},\quad F\circ\sigma=\mu_{i-1}^{-1}\circ F\circ\mu_{i}\ (1<i<l),\quad F\circ\sigma=\mu_{l-1}^{-1}\circ F.$
Set $\mu_{0}(z)=z$ when $l=1$. Then $\tau:=\mu_{l-1}^{-1}$ satisfies $F\circ\sigma=\tau\circ F$. We claim $\tau\in E_{0}(F)$. For $l=1$, this is clear. For $l\geq 2$, note that
$F^{\circ l}=(F\circ\sigma)^{\circ l}=(\tau\circ F)^{\circ l}.$
By [Pak25a, Lemma 3.5], $\tau\in\mathrm{Aut}(F^{\circ l})\subseteq\mathrm{Aut}_{\infty}(F)\subseteq E_{0}(F)$. By the maximality in the definition of $G_{0}(F)$, we obtain $E_{0}(F)\subseteq G_{0}(F)$.
Thus, $G_{0}(F)=\mathrm{Aut}(F^{\circ s})=\mathrm{Aut}_{\infty}(F)=E_{0}(F)\subseteq\mathrm{Aff}(\mathbb{C})$.